# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

### v1

@@ 5 Conclusion and Discussion
Under OLS, the $\bm{\psi}$} of the linear regression for each subject:
${\bm{\psi}}^{(i)}=X_{i}(y_{i}-X_{i}^{\top}{\bm{\theta}}),$
and the expected first-order derivative of ${\bm{\psi}}^{(i)}$:
$\dot{{\bm{\psi}}}=\mathbb{E}\left[X_{i}X_{i}^{\top}\right].$
Similarly we can obtain ${\bm{\psi}}^{(i)}_{k}$ and $\dot{{\bm{\psi}}}_{k}$ by substituting the $y_{i}$ and $X_{i}$ with their imputed counterpart under the $k$-th missingness pattern. As such, taking $\Sigma_{\bm{\theta}}$ for an example, denoting $N_{\textrm{CC}}=\sum^{N}_{i=1}I({\mathcal{C}}_{i}=\infty)$, we can compute the consistent estimate of the meat:
$\widehat{\mathbb{E}}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]=\frac{1}{N_{\textrm{CC}}}\sum^{N}_{i=1}I({\mathcal{C}}_{i}=\infty)\frac{X_{i}(y_{i}-X_{i}^{\top}\widehat{{\bm{\theta}}}_{\text{WCC}})^{2}X^{\top}_{i}}{\widehat{\pi}_{\infty}(X_{i},y_{i})}.$
For the bread:
$\widehat{\dot{{\bm{\psi}}}}=\frac{1}{N_{\textrm{CC}}}\sum^{N}_{i=1}I({\mathcal{C}}_{i}=\infty)X_{i}X_{i}^{\top}.$
Note that $\widehat{{\bm{\theta}}}_{\text{WCC}}$ is plugged in to compute ${\bm{\psi}}^{(i)}$. For the corresponding meat and bread parts in all the other variance-covariance matrices, we could apply the similar procedure to obtain their consistent estimates. When computing ${\bm{\psi}}_{k}$ and $\dot{{\bm{\psi}}}_{k}$, $\widehat{{\bm{\gamma}}}_{1,k}$ will be plugged in into the estimating equations.
@@
In the logistic regression, the procedure for obtaining the consistent estimates of the variance-covariance matrices are similar; the only differences are the specific form of ${\bm{\psi}}^{(i)}$, ${\bm{\psi}}^{(i)}_{k}$, $\dot{{\bm{\psi}}}$, and $\dot{{\bm{\psi}}}_{k}$. Specifically, under the logistic regression setting, the log-likelihood for each subject is:
@@
${\bm{\psi}}^{(i)}=X_{i}\left(y_{i}-\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}\right),$
and
$\dot{{\bm{\psi}}}=\mathbb{E}\left[X_{i}X^{\top}_{i}\mu_{i}(1-\mu_{i})\right],$
where $\mu_{i}=\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}$. We can similarly obtain ${\bm{\psi}}^{(i)}_{k}$ and $\dot{{\bm{\psi}}}_{k}$ by substituting the $y_{i}$ and $X_{i}$ with their imputed counterpart under the $k$-th missingness pattern.
@@
In this subsection we make comment on why $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is an AIPW estimator. Specifically, when the propensity score model is known, the $i$-th influence function of the AIPW estimators, as defined in Tsiatis, (2006), is:
$\varphi_{\text{Tsiatis},i}$ $=-\mathbb{E}\left\{\frac{\partial{\bm{\psi}}}{\partial{\bm{\theta}}}\Bigr{|}_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right\}^{-1}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})+\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\left(\sum^{K}_{k=1}\pi_{k}(G_{k}(\bm{Y}_{i}))L_{2k}(G_{k}(\bm{Y}_{i}))\right)\right$.
@@
$=-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{(i)}-\sum^{K}_{k=1}\left(\frac{I({\mathcal{C}}_{i}=\infty)\pi_{k}(G_{k}(\bm{Y}_{i}))}{\pi_{\infty}(\bm{Y}_{i})}-I({\mathcal{C}}=k)\right)\dot{{\bm{\psi}}}^{-1}L_{2k}(G_{k}(\bm{Y}_{i})).$
where $L_{2k}(G_{k}(\bm{Y}_{i}))$ are a series of arbitrarily chosen functions. By carefully designing $L_{2k}(G_{k}(\bm{Y}_{i}))$ as:
$L_{2k}(G_{k}(\bm{Y}_{i}))=-\frac{1}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\dot{{\bm{\psi}}}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}{\bm{\psi}}^{(i)}_{k}$,
we have $\varphi_{\text{Tsiatis},i}={\bm{\Gamma}}_{i}$, which suggests that the $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ estimator is an AIPW estimator when the propensity score model is known. Similar conclusion could be derived when the propensity score model is unknown but correctly specified, as one only need to further expand $\widehat{{\bm{\omega}}}$ in ${\bm{\Gamma}}_{i}$ (and $\varphi_{\text{Tsiatis},i}$) to obtain $\varphi_{i}$.
@@
Solving the equation we know that
@@
is positive definite matrix. As a result, we have $\Sigma_{\text{PS-PPI}}\preceq\Sigma_{{\bm{\theta}}}$.
@@
From Gronsbell et al., (2024), we know that
$\textbf{IF}_{i,\text{Chen}}=-\dot{{\bm{\psi}}}^{-1}\left\{{\bm{\psi}}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}+\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\text{Var}({\bm{\psi}}_{1})^{-1}{\bm{\psi}}_{1}\frac{\pi_{\infty}-I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}\right\}.$
Recall that when $\pi(\cdot)$ is known, the $i$-th influence function of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is:
$\textbf{IF}_{i,\text{PS-PPI}}=-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}_{k}.$
@@
$\textbf{IF}_{i,\text{PS-PPI}}=-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}+\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}-\frac{I({\mathcal{C}}_{i}=1)}{\pi_{1}}\right){\bm{\psi}}_{1}.$
@@
$\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,1}}=\dot{{\bm{\psi}}}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}_{1}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty},$
$\Sigma_{{\bm{\gamma}}_{1,1}}=\dot{{\bm{\psi}}}^{-1}_{1}\mathbb{E}\left[\frac{{\bm{\psi}}_{1}{\bm{\psi}}^{\top}_{1}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty},$
$\Sigma_{{\bm{\gamma}}_{2,1}}=\dot{{\bm{\psi}}}^{-1}_{1}\mathbb{E}\left[\frac{{\bm{\psi}}_{1}{\bm{\psi}}^{\top}_{1}}{\pi_{1}(G_{1}(\bm{Y}_{i}))}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}(1-\pi_{\infty})^{-1}.$
@@
$\mathbf{W}^{\ast}$ $=\left(\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty}\right)\times\left(\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\right)^{-1}\pi_{\infty}\pi_{1}$
$=\left(\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty}\right)\times\left(\dot{{\bm{\psi}}}^{\top}_{1}\text{Var}({\bm{\psi}}_{1})^{-1}\dot{{\bm{\psi}}}_{1}\right)\pi_{\infty}\pi_{1}$
$=\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\text{Var}({\bm{\psi}}_{1})^{-1}\dot{{\bm{\psi}}}_{1}\pi_{1}.$
@@
$\textbf{IF}_{i,\text{PS-PPI}}$ $=-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}{\bm{\psi}}+\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\text{Var}({\bm{\psi}}_{1})^{-1}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}-\frac{I({\mathcal{C}}_{i}=1)}{\pi_{1}}\right){\bm{\psi}}_{1}\pi_{1}$
$=-\dot{{\bm{\psi}}}^{-1}\left\{{\bm{\psi}}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}+\text{Cov}({\bm{\psi}},{\bm{\psi}}_{1})\text{Var}({\bm{\psi}}_{1})^{-1}{\bm{\psi}}_{1}\frac{\pi_{\infty}-I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}\right\}=\textbf{IF}_{i,\text{Chen}}.$
@@
From Kluger et al., (2025), we know that the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PTD}}$ is
@@
In this subsection we present the additional experimental results of the synthetic data simulation under different simulation settings. Specifically, instead of fixing the prediction bias level, we fix the prediction noise level at $\sigma_{\text{pred}}=0$ and vary the bias level and present the results in Figure S1. Based on the simulation settings adopted in Figure 2 and S1, we correspondingly substitute the estimated propensity score with the real propensity score and present the simulation results in Figure S2 and S3. Similarly, we substitute the estimated propensity score with the one obtained from the misspecified model and present the simulation resutls in Figure S4 and S5.
Figure S1: The same as Figure 2, except that the noise level is fixed at $\sigma_{\text{pred}}=0$, and X-axis in all panels represents the prediction bias level used in the simulations.

### v2

@@ Acknowledgment
The $\bm{\psi}$} of the linear regression for each subject:
${\bm{\psi}}^{(i)}=X_{i}(y_{i}-X_{i}^{\top}{\bm{\theta}}).$
From Eq. (S1.3), the meat of $\Sigma_{{\bm{\theta}}}$ is
$\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]=\mathbb{E}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}^{2}(X_{i},y_{i})}X_{i}(y_{i}-X^{\top}_{i}{\bm{\theta}}^{\ast})^{2}X^{\top}_{i}\right].$
On the other hand, the bread of $\Sigma_{{\bm{\theta}}}$ is:
$\dot{{\bm{\psi}}}_{\infty}=\mathbb{E}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi(X_{i},y_{i})}X_{i}X_{i}^{\top}\right].$
Therefore, we can compute the consistent estimate of $\Sigma_{\bm{\theta}}$ by estimating the bread and the meat using data:
$\widehat{\mathbb{E}}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}^{2}_{\infty}(X_{i},y_{i})}X_{i}(y_{i}-X_{i}^{\top}\widehat{{\bm{\theta}}}_{\text{WCC}})^{2}X^{\top}_{i}=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)\widehat{\epsilon}^{2}_{i}}{\widehat{\pi}^{2}_{\infty}(X_{i},y_{i})}X_{i}X^{\top}_{i},$
$\widehat{\mathbb{E}}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi(X_{i},y_{i})}X_{i}X_{i}^{\top}\right]=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(X_{i},y_{i})}X_{i}X^{\top}_{i},$
where $\widehat{\epsilon}_{i}=y_{i}-X_{i}^{\top}\widehat{{\bm{\theta}}}_{\text{WCC}}$. For the corresponding meat and bread parts in all the other variance-covariance matrices, we could apply a similar procedure to obtain their consistent estimates.
@@
In the logistic regression, the procedure for obtaining the consistent estimates of the variance-covariance matrices are similar; the only difference is the closed-form expression of ${\bm{\psi}}^{\ast(i)}$, $\dot{{\bm{\psi}}}_{\infty}$, ${\bm{\psi}}^{\ast(i)}_{k}$, and $\dot{{\bm{\psi}}}_{k}$. Specifically, under the logistic regression setting, the log-likelihood for each subject is:
@@
${\bm{\psi}}^{(i)}=X_{i}\left(y_{i}-\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}\right).$
As such, the meat is:
$\widehat{\mathbb{E}}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}^{2}_{\infty}(X_{i},y_{i})}X_{i}\left(y_{i}-\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}\right)^{2}X^{\top}_{i}.$
Similarly, since $\dot{{\bm{\psi}}}_{\infty}=\mathbb{E}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi(X_{i},y_{i})}X_{i}X^{\top}_{i}\mu_{i}(1-\mu_{i})\right]$, where $\mu_{i}=\frac{e^{X_{i}^{\top}{\bm{\theta}}}}{1+e^{X_{i}^{\top}{\bm{\theta}}}}$, the bread part is:
$\widehat{\mathbb{E}}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi(X_{i},y_{i})}X_{i}X^{\top}_{i}\mu_{i}(1-\mu_{i})\right]=\frac{1}{N}\sum^{N}_{i=1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(X_{i},y_{i})}X_{i}X^{\top}_{i}\mu_{i}(1-\mu_{i}).$
For the corresponding meat and bread parts in all the other variance-covariance matrices, we could apply a similar procedure to obtain their consistent estimates.
@@
In this subsection, we make a comment on why $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is an AIPW estimator. Specifically, when the propensity score model is known, the $i$-th influence function of the AIPW estimators, as defined in tsiatis2006semiparametric, is:
$\varphi_{\text{Tsiatis},i}$ $=-\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]^{-1}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})+\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\left(\sum^{K}_{k=1}\pi_{k}(G_{k}(\bm{Y}_{i}))L_{2k}(G_{k}(\bm{Y}_{i}))\right)\right$.
@@
$=-\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{\ast(i)}-\sum^{K}_{k=1}\left(\frac{I({\mathcal{C}}_{i}=\infty)\pi_{k}(G_{k}(\bm{Y}_{i}))}{\pi_{\infty}(\bm{Y}_{i})}-I({\mathcal{C}}=k)\right)\dot{{\bm{\psi}}}_{\infty}^{-1}L_{2k}(G_{k}(\bm{Y}_{i})).$
Note that $\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]^{-1}=-\dot{{\bm{\psi}}}_{\infty}^{-1}$ holds when the propensity score is known:
$\dot{{\bm{\psi}}}_{\infty}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]=\mathbb{E}\left[\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\frac{\partial{\bm{\psi}}^{(i)}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right].$
$L_{2k}(G_{k}(\bm{Y}_{i}))$ are a series of arbitrarily chosen functions. By carefully designing $L_{2k}(G_{k}(\bm{Y}_{i}))$ as:
$L_{2k}(G_{k}(\bm{Y}_{i}))=-\frac{1}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\dot{{\bm{\psi}}}_{\infty}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}{\bm{\psi}}^{\ast(i)}_{k}$,
we have $\varphi_{\text{Tsiatis},i}={\bm{\Gamma}}^{(i)}$, which suggests that the $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ estimator is an AIPW estimator when the propensity score model is known. A similar conclusion could be derived when the propensity score model is unknown but correctly specified, as one only needs to further expand $\widehat{{\bm{\omega}}}$ in ${\bm{\Gamma}}^{(i)}$ (and $\varphi_{\text{Tsiatis},i}$) to obtain the $i$-th influence function of the PS-PPI estimator $\varphi_{i}$.
@@
Solving the equation, we know that
@@
is a positive definite matrix. As a result, we have $\Sigma_{\text{PS-PPI}}\preceq\Sigma_{{\bm{\theta}}}$.
@@
From gronsbell2024another, we know that
$\textbf{IF}_{i,\text{Chen}}=-\dot{{\bm{\psi}}}_{\infty}^{-1}\left\{{\bm{\psi}}^{\ast(i)}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}+\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}{\bm{\psi}}^{\ast(i)}_{1}\frac{\pi_{\infty}-I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}\right\}.$
Note that $\dot{{\bm{\psi}}}_{\infty}^{-1}$ introduced in the above equation is equivalent to the $\mathbf{A}=\mathbb{E}\left[\mathbf{X}\mathbf{X}^{\top}\right]^{-1}$ introduced in gronsbell2024another in the linear regression case they considered. On the other hand, recall that when $\pi(\cdot)$ is known, the $i$-th influence function of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is:
$\textbf{IF}_{i,\text{PS-PPI}}=-\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{\ast(i)}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}^{\ast(i)}_{k}$.
@@
$\textbf{IF}_{i,\text{PS-PPI}}=-\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}{\bm{\psi}}^{\ast(i)}+\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}-\frac{I({\mathcal{C}}_{i}=1)}{\pi_{1}}\right){\bm{\psi}}^{\ast(i)}_{1}$.
@@
$\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,1}}=\dot{{\bm{\psi}}}_{\infty}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}_{1}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}_{\infty}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty},$
$\Sigma_{{\bm{\gamma}}_{1,1}}=\dot{{\bm{\psi}}}^{-1}_{1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}_{1}{\bm{\psi}}^{\ast(i)\top}_{1}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty},$
$\Sigma_{{\bm{\gamma}}_{2,1}}=\dot{{\bm{\psi}}}^{-1}_{1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}_{1}{\bm{\psi}}^{\ast(i)\top}_{1}}{\pi_{1}(G_{1}(\bm{Y}_{i}))}\right]\dot{{\bm{\psi}}}^{-1\top}_{1}=\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}(1-\pi_{\infty})^{-1}.$
@@
$\mathbf{W}^{\ast}$ $=\left(\dot{{\bm{\psi}}}_{\infty}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty}\right)\times\left(\dot{{\bm{\psi}}}^{-1}_{1}\text{Var}({\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\right)^{-1}\pi_{\infty}\pi_{1}$
$=\left(\dot{{\bm{\psi}}}_{\infty}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\dot{{\bm{\psi}}}^{-1\top}_{1}\pi^{-1}_{\infty}\right)\times\left(\dot{{\bm{\psi}}}^{\top}_{1}\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}\dot{{\bm{\psi}}}_{1}\right)\pi_{\infty}\pi_{1}$
$=\dot{{\bm{\psi}}}_{\infty}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}\dot{{\bm{\psi}}}_{1}\pi_{1}.$
@@
$\textbf{IF}_{i,\text{PS-PPI}}$ $=-\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}{\bm{\psi}}^{\ast(i)}$
$+\dot{{\bm{\psi}}}^{-1}\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}-\frac{I({\mathcal{C}}_{i}=1)}{\pi_{1}}\right){\bm{\psi}}^{\ast(i)}_{1}\pi_{1}$
$=-\dot{{\bm{\psi}}}_{\infty}^{-1}\left\{{\bm{\psi}}^{\ast(i)}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}+\text{Cov}({\bm{\psi}}^{\ast(i)},{\bm{\psi}}^{\ast(i)}_{1})\text{Var}({\bm{\psi}}^{\ast(i)}_{1})^{-1}{\bm{\psi}}^{\ast(i)}_{1}\frac{\pi_{\infty}-I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}}\right\}$
$=\textbf{IF}_{i,\text{Chen}}.$
@@
From kluger2025prediction, we know that the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PTD}}$ is
@@
In this subsection, we present the additional experimental results of the synthetic data simulation under different simulation settings. Specifically, instead of fixing the prediction bias level, we fix the prediction noise level at $\sigma_{\text{pred}}=0$ and vary the bias level and present the results in Figure S1. Based on the simulation settings adopted in Figure 2 and S1, we correspondingly substitute the estimated propensity score with the real propensity score and present the simulation results in Figure S2 and S3. Similarly, we substitute the estimated propensity score with the one obtained from the misspecified model and present the simulation results in Figure S4 and S5.
Figure S1: The same as Figure 2, except that the noise level is fixed at $\sigma_{\text{pred}}=0$, and the X-axis in all panels represents the prediction bias level used in the simulations.